# NSmark: Null Space Based Black-box Watermarking Defense Framework for Pre-trained Language Models

### v1

@@ 5 Further Analysis
BERT NSMD $2.94\times 10^{-6}$ 29.77 25.29 22.52 21.96 24.37
@@
RoBERTa NSMD $2.53\times 10^{-6}$ 50.17 30.97 25.15 26.78 28.43
@@
DeBERTa NSMD $2.91\times 10^{-6}$ 31.89 25.74 23.52 27.23 37.69
@@
XLNet NSMD $2.90\times 10^{-6}$ 24.12 23.52 25.29 24.06 26.20
@@
Table 7: Impact of using a validation set without triggers on watermarking performance.
Next we discuss the necessity of using the verification set with trigger to calculate NSMD. In §$\S$ 3.4.1 of watermark generation, we introduce the selection and generation method of the verification trigger set. In order to explain why uses the trigger set instead of the clean set to form $D_{V}$, we compare the effects of trigger set (NSMD) with clean set ($\text{NSMD}_{c}$) in Table 7. It shows that for $f_{wm}$, NSMD is always close to 00, but the model shows a huge difference after downstream fine-tuning. In this case, $\text{NSMD}_{c}$ is significantly higher than NSMD, indicating that the output matrix no longer matches the original null space matrix. This is because only the output of triggered set is mapped to $sig$ through $E$. Results in Table 5 show that fine-tuning has little effect on WER, indicating that fine-tuning has little effect on the output representation of trigger set. This means the output matrix of trigger set changes little, so it still matches original $N$, and NSMD is always close to 00. However, after fine-tuning, the output representation corresponding to the clean data set will change significantly to achieve better performance on different downstream tasks, which causes the output matrix no longer match the original $N$, causing $\text{NSMD}_{c}$ to significantly increase. Therefore, the trigger set is needed for verifying null space.
@@ 6 Conclusion
This paper proposes NSmark, a black-box watermark framework for ownership verification using the output of PLM. We first analyze and introduce LL-LFEA, and propose a solution that can use null space invariance for watermark verification. We conduct an overall design from three aspects: watermark generation, watermark embedding, and watermark verification. Two indicators, WER and NSMD, are used to jointly verify the existence and identity of the watermark. Experiments prove the effectiveness, reliability and fidelity of NSmark, and it has satisfactory performance under various attacks. With the cooperation of two verification methods, it works a secure and robust watermarking scheme.
7 Limitations
Our approach has limitations in two main aspects. First, our method only focuses on LMs as we use the vocab to build connection between $sig$ and trigger. Solutions applicable to other models (such as visual or multi-modal models) need further exploration. Second, further research is needed on this basis to further protect the downstream network.
@@ Appendix A Additional Details of Theory and Algorithm
A.1 Proofs of Null Space Verification Theory
@@
The null space $N(A)$ of matrix $A_{(a\times b)}$ is the set of all $b$-dimensional vectors $x$ that satisfy $Ax=\vec{0}$ Axler (2015). That is, $N(A)=\{x\in\mathbb{R}^{b},Ax=\vec{0}\}$. Using $f_{wm}$ to denote PLM embedded with watermark, assuming $A_{1(n\times m)}=\{f_{wm}(x),x\in D_{T}\}$ is the matrix concatenated from the output vectors, where $D_{T}$ is the verification dataset with watermark trigger, $m$ is the size of $D_{T}$ and $n$ is the dimension of the output vector of the last layer of PLM. Let the null space matrix of $A_{1}$ be $N_{1}$, then $A_{1}\times N_{1}=\textbf{0}$.
After performing LL-LFEA, assuming the new output matrix of $f_{wm}(D_{T})$ is $A_{2(n\times m)}$, according to Sec.3.1, we have $A_{2}=Q\times A_{1}$. Then $A_{2}\times N_{1}=(Q\times A_{1})\times N_{1}=Q\times(A_{1}\times N_{1})=\textbf{0}$, which means $N_{1}$ belongs to the null space matrix of $A_{2}$. As $Q$ is a reversible matrix, then $rank(A_{1})=rank(A_{2})$, and the null space of $A_{1}$ and $A_{2}$ have the same dimension. It can be concluded that $N_{1}$ is also the null space matrix of $A_{2}$. ∎
A.2 Estimation of NSMD
We define NSMD by introducing the distribution of elements in a matrix, which is obtained by matrix multiplication of the output matrix $A$ of any PLM without watermark and the null space matrix $N$ of $f_{wm}$. In $H_{(n\times p)}=A_{(n\times m)}\times N_{(m\times p)}$, $H_{j,j}=\alpha_{i}\cdot\beta_{j}$ is the dot product of the i-th row vector of $A$ and the j-th column vector of $N$. It is proposed that the approximate distribution of the angle between $n$ random uniformly distributed unit vectors in space $\mathbb{R}^{m}$ Cai et al. (2013). In space $\mathbb{R}^{m}$, given two random vectors uniformly distributed on the unit sphere, the angle $\theta$ between the two random vectors converges to a distribution whose probability density function is:
@@ A.2 Estimation of NSMD
When $m=2$, $f(\theta)$ is uniformly distributed on $[0,\pi]$; when $m>2$, $f(\theta)$ has a single peak at $\theta=\frac{\pi}{2}$. When $m>5$, the distribution of $f(\theta)$ is very close to the normal distribution. Most of the $C_{m}^{2}$ angles formed by $m$ randomly uniformly distributed unit vectors are concentrated around $\frac{\pi}{2}$, and this clustering will enhance with the increase of the dimension $m$, because if $\theta\neq\frac{\pi}{2}$, then $(\sin\theta)^{m-2}$ will converge to 00 faster. This shows that in high-dimensional space, two randomly selected vectors are almost orthogonal.
We further derive the distribution of the dot product of two random vectors uniformly distributed and independently selected on the unit ball in space $\mathbb{R}^{m}$. Let $\alpha$ and $\beta$ be unit vectors and $\theta$ be the angle between them, then $\alpha\cdot\beta=\cos(\theta)$. It is known that $\theta$ obeys the probability distribution $f(\theta)$, then the probability density function of $y=\alpha\cdot\beta=\cos(\theta),y\in[-1,1]$ is:
@@
As $m$ increases, $DY$ gradually approaches 00. Figure 8 shows the relationship between the DY and the $m$, and Table 8 shows the specific values of the variance when $m$ takes specific values. Under the common output dimension of PLM, that is, when $m=1000$ or so, $DY$ is still a distance to 00.
@@
A.3 Trigger Generation Algorithm
@@ A.4 Select Algorithm
A.5 Process of Spread Spectrum Modulation and Despread Spectrum
A.5.1 Spread Spectrum Modulation
Assume that the digital signature $sig$ is $n$ bits, $sig=\{a_{i}|a_{i}\in\{-1,1\},i\in[0,n-1]\}$, and set the spreading factor to $k$. Expand $sig$ horizontally by $k$ times to obtain $sig_{repeat}=\{ra_{j}|ra_{j}=a_{i},i=j\ mod\ n,ra_{j}\in\{-1,1\},j\in[0,k\times n-1]\}$. Input $sig$ as a seed into the pseudo-random generator to obtain the key $sm=\{b_{j}|b_{j}\in\{-1,1\},j\in[0,k\times n-1]\}$ for spread spectrum modulation. Use $sm$ to modulate $sig_{repeat}$ to obtain the spread spectrum modulated digital signature $sig_{sm}=\{sa_{j}|sa_{j}=ra_{j}\times b_{j},j\in[0,k\times n-1]\}$. Figure 4 shows an example of 3×3$\times$3 × spreading.
A.5.2 Despread Spectrum
Let the mapping vector output by $E$ be $O=\{o_{j},j\in[0,k\times n-1]\}$, modulate it with $sm=\{b_{j}|b_{j}\in\{-1,1\},j\in[0,k\times n-1]\}$ to get $O_{repeat}=\{ro_{j}|ro_{j}=o_{j}/b_{j},j\in[0,k\times n-1]\}$, then quantify $O_{repeat}$ to get $O_{quan}=\{qo_{j}|qo_{j}\in\{-1,0,1\},j\in[0,k\times n-1]\}$. At last the signature is extracted by counting the number of $n$ positions that appears most often in $k$ copies. $sig^{\prime}=\{a^{\prime}_{i}|a^{\prime}_{i}\in\{-1,0,1\},i\in[0,n-1]\}$. The quantification method is shown as follows:
@@ Appendix B Additional Experimental Results and Analyses
B.1 Defense against LL-LFEA+finetuning
@@
Table 9: Impact of LL-LFEA+ fine-tuning attack on watermark performance.
After applying LL-LFEA, the attacker may add a network to $f_{LL-LFEA}$ and fine-tune it for downstream tasks. We hope the model after the LL-LFEA+fine-tuning attack can still maintain the watermark. Table 9 shows results on different PLMs and different downstream tasks are not exactly the same. WER of different models has decreased significantly to varying degrees. Most NSMDs are still below the threshold, but RoBERTa and DeBERTa change more on SST-2, which is generally consistent with that of fine-tuning without LL-LFEA attack (Table 5). Through ACC, we can find that LL-LFEA attack does not affect the performance of the model on the original task.
B.2 Robustness against Fine-pruning
@@
Table 10: Impact of fine-pruning attack on watermark performance.
Figure 9: The t-SNE visualization of output feature vectors of watermarked models. (1) Left column: $f_{wm}$ on WikiText; (2) Middle column: $F_{wm}$ on SST-2; (3) Right column: $F_{wm}$ on SST-5.
B.3 Feature Visualization
To demonstrate the effectiveness of our scheme, we use t-SNE to visualize the feature distribution of the watermarking model. As shown in Figure 9, the input with trigger and the input without trigger can be well separated in the output of PLM and $E$, whether for $f_{wm}$ or fine-tuned $F_{wm}$.

### v2

@@ 5 Further Analysis
Table 7: The necessity of using trigger set in verification. $\text{NSMD}_{t}$ is the result for trigger set and $\text{NSMD}_{c}$ is for clean set.
@@
BERT $\text{NSMD}_{t}$ $2.94\times 10^{-6}$ 29.77 25.29 22.52 21.96 24.37
@@
RoBERTa $\text{NSMD}_{t}$ $2.53\times 10^{-6}$ 50.17 30.97 25.15 26.78 28.43
@@
DeBERTa $\text{NSMD}_{t}$ $2.91\times 10^{-6}$ 31.89 25.74 23.52 27.23 37.69
@@
XLNet $\text{NSMD}_{t}$ $2.90\times 10^{-6}$ 24.12 23.52 25.29 24.06 26.20
@@
Next we discuss the necessity of using trigger set in verification rather than clean set. As shown in Table 7, before downstream fine-tuning $(f_{wm})$, NSMD for trigger set ($\text{NSMD}_{t}$) and clean set ($\text{NSMD}_{c}$) are all close to 00.After downstream fine-tuning, $\text{NSMD}_{c}$ is significantly higher than $\text{NSMD}_{t}$.
Combining these results, we think fine-tuning has little effect on the output representation of trigger set. However, the output representation of the clean set will change significantly for better performance on different downstream tasks, which causes the null space matrix no longer match the original $N$. Thus $\text{NSMD}_{c}$ has significantly increase. Therefore, the trigger set is needed for verifying null space.
@@ 6 Conclusion
This paper proposes NSmark, a black-box watermark framework for verification of ownership using the output of LMs. We first analyze and introduce LL-LFEA, and propose a solution that can use null space invariance for watermark verification. We conduct an overall design from three aspects: watermark generation, watermark embedding, and watermark verification. Two indicators, WER and NSMD, are used to jointly verify the existence and identity of the watermark. Experiments demonstrate the effectiveness, scalability, reliability, and fidelity of NSmark, and it has satisfactory performance under various attacks. With the cooperation of two verification methods, a robust and secure watermarking scheme works.
Impact Statement
LM watermarking plays a crucial role in IP protection, with its significant societal implications. This paper highlights the vulnerabilities of existing watermarking schemes, particularly in the context of LL-LFEA threats, emphasizing the urgent need for further research and practical deployment of robust black-box LM watermarks. We believe there is no direct negative impact of making our findings public, nor is there a clear avenue for responsible disclosure. Conversely, we assert that our work contributes positively to society by exposing the vulnerabilities in current watermarking schemes, underscoring the necessity for more resilient designs and rigorous evaluation methodologies. This research represents a significant step toward the practical implementation of robust LM watermarks.
@@ Appendix A Additional Details of Theory and Algorithm
A.1 The generation of $Q$
In principle, $Q$ needs to be a reversible matrix. To make the value of $Q$ more stable for tensor calculations, we choose to constrain each element to be uniformly distributed between $[0,1]$. The method to generate such a $Q$ is very simple, just sample uniformly between $[0,1]$, and the resulting matrix is a reversible matrix (probability very close to 100%). The probability that the random matrix $Q$ is singular is the same as a point in $R^{n^{2}}$ lands in the zero set of a polynomial, which has Lebesgue measure 00 (Fleming, 2012).
A.2 Proofs of Null Space Verification Theory
@@
The null space $N(A)$ of the matrix $A_{(a\times b)}$ is the set of all $b$-dimensional vectors $x$ that satisfy $Ax=\vec{0}$ (Axler, 2015). That is, $N(A)=\{x\in\mathbb{R}^{b},Ax=\vec{0}\}$. Using $f_{wm}$ to denote the LM embedded with the watermark, assuming $A_{1(n\times m)}=\{f_{wm}(x),x\in D_{T}\}$ is the matrix concatenated from the output vectors, where $D_{T}$ is the verification dataset with watermark trigger, $m$ is the size of $D_{T}$ and $n$ is the dimension of the output vector of the last layer of the LM. Let the null space matrix of $A_{1}$ be $N_{1}$, then $A_{1}\times N_{1}=\textbf{0}$.
After performing LL-LFEA, assuming that the new output matrix of $f_{wm}(D_{T})$ is $A_{2(n\times m)}$, according to Section 3.1, we have $A_{2}=Q\times A_{1}$. Then $A_{2}\times N_{1}=(Q\times A_{1})\times N_{1}=Q\times(A_{1}\times N_{1})=\textbf{0}$, which means that $N_{1}$ belongs to the null space matrix of $A_{2}$. As $Q$ is a reversible matrix, then $rank(A_{1})=rank(A_{2})$, and the null spaces of $A_{1}$ and $A_{2}$ have the same dimension. It can be concluded that $N_{1}$ is also the null space matrix of $A_{2}$. ∎
A.3 Estimation of NSMD
We define NSMD by introducing the distribution of elements in a matrix, which is obtained by matrix multiplication of the output matrix $A$ of any LM without watermark and the null space matrix $N$ of $f_{wm}$. In $H_{(n\times p)}=A_{(n\times m)}\times N_{(m\times p)}$, $H_{j,j}=\alpha_{i}\cdot\beta_{j}$ is the dot product of the i-th row vector of $A$ and the j-th column vector of $N$. The approximate distribution of the angle between $n$ random uniformly distributed unit vectors in space $\mathbb{R}^{m}$ (Cai et al., 2013). In space $\mathbb{R}^{m}$, given two random vectors uniformly distributed on the unit sphere, the angle $\theta$ between the two random vectors converges to a distribution whose probability density function is:
@@ A.3 Estimation of NSMD
When $m=2$, $f(\theta)$ is uniformly distributed on $[0,\pi]$; when $m>2$, $f(\theta)$ has a single peak at $\theta=\frac{\pi}{2}$. When $m>5$, the distribution of $f(\theta)$ is very close to the normal distribution. Most of the $C_{m}^{2}$ angles formed by $m$ unit vectors randomly uniformly distributed are concentrated around $\frac{\pi}{2}$, and this clustering will enhance with the increase of the dimension $m$, because if $\theta\neq\frac{\pi}{2}$, then $(\sin\theta)^{m-2}$ will converge to 00 faster. This shows that in high-dimensional space, two randomly selected vectors are almost orthogonal.
We further derive the distribution of the dot product of two random vectors uniformly distributed and independently selected on the unit ball in space $\mathbb{R}^{m}$. Let $\alpha$ and $\beta$ be unit vectors and let $\theta$ be the angle between them, then $\alpha\cdot\beta=\cos(\theta)$. It is known that $\theta$ obeys the probability distribution $f(\theta)$, then the probability density function of $y=\alpha\cdot\beta=\cos(\theta),y\in[-1,1]$ is:
@@
As $m$ increases, $DY$ gradually approaches 00. Figure 8 shows the relationship between the DY and the $m$, and Table 8 shows the specific values of the variance when $m$ takes specific values. Under the common output dimension of LM, that is, when $m=1000$ or so, $DY$ is still a distance to 00.
@@
A.4 Trigger Generation Algorithm
@@ A.5 Select Algorithm
A.6 Process of Spread Spectrum Modulation and Despread Spectrum
A.6.1 Spread Spectrum Modulation
Assume that the digital signature $sig$ is $n$ bits, $sig=\{a_{i}|a_{i}\in\{-1,1\},i\in[0,n-1]\}$, and set the spread factor to $k$. Expand $sig$ horizontally by $k$ times to obtain $sig_{repeat}=\{ra_{j}|ra_{j}=a_{i},i=j\ mod\ n,ra_{j}\in\{-1,1\},j\in[0,k\times n-1]\}$. Input $sig$ as a seed in the pseudo-random generator to obtain the key $sm=\{b_{j}|b_{j}\in\{-1,1\},j\in[0,k\times n-1]\}$ for spread spectrum modulation. Use $sm$ to modulate $sig_{repeat}$ to obtain the spread spectrum modulated digital signature $sig_{sm}=\{sa_{j}|sa_{j}=ra_{j}\times b_{j},j\in[0,k\times n-1]\}$. Figure 4 shows an example of 3×3$\times$3 × spreading.
A.6.2 Despread Spectrum
Despread spectrum is the inverse process of the spread spectrum (detailed process in Appendix A.6.2). Let the output of the mapping vector by $E$ be $O=\{o_{j},j\in[0,k\times n-1]\}$, modulate it with $sm=\{b_{j}|b_{j}\in\{-1,1\},j\in[0,k\times n-1]\}$ to get $O_{repeat}=\{ro_{j}|ro_{j}=o_{j}/b_{j},j\in[0,k\times n-1]\}$, then quantify $O_{repeat}$ to get $O_{quan}=\{qo_{j}|qo_{j}\in\{-1,0,1\},j\in[0,k\times n-1]\}$. Finally, the signature is extracted by counting the number of $n$ positions that appear most often in $k$ copies. $sig^{\prime}=\{a^{\prime}_{i}|a^{\prime}_{i}\in\{-1,0,1\},i\in[0,n-1]\}$. The quantification method is shown as follows:
@@ A.6.2 Despread Spectrum
At last the signature is extracted as $sig^{\prime}=\{a^{\prime}_{i}|a^{\prime}_{i}\in\{-1,0,1\},i\in[0,n-1]\}$.
@@ Appendix B Additional Experimental Results and Analyses
B.1 LL-LFEA attack results on EmbMarker
As shown in Table 9, as defined in (Peng et al., 2023), the difference in cosine similarity ($\Delta_{cos}$), the difference of squared L2 distance ($\Delta_{l2}$), and the p-value of the KS test are used to measure the effectiveness of watermark. RedAlarm (Zhang et al., 2023) is another attack baseline work that demonstrates the effectiveness of EmbMarker. After the LL-LFEA attack, all the metrics of EmbMarker are very close to those of the original and RedAlarm, which fully demonstrates the effectiveness of LL-LFEA on existing watermarking schemes.
Table 9: Results of LL-LFEA attack on EmbMarker. For watermark, $\uparrow$ means higher metrics are better. $\downarrow$ means lower metrics are better. In contrast, after LL-LFEA attack, the higher p-value, $\Delta_{l2}\%$ and lower $\Delta_{cos}\%$ compared to EmbMarker can illustrate the effectiveness of LL-LFEA.
Dataset Method ACC(%) p-value $\downarrow$ $\Delta_{cos}\%\uparrow$ $\Delta_{l2}\%\downarrow$
SST2 Original 93.76 >0.34 -0.07 0.14
RedAlarm 93.76 >0.09 1.35 -2.70
EmbMarker 93.55 <$10^{-5}$ 4.07 -8.13
EmbMarker+LL-LFEA 92.43 0.01 0.14 -0.28
MIND Original 77.30 >0.08 -0.76 1.52
RedAlarm 77.18 >0.38 -2.08 4.17
EmbMarker 77.29 <$10^{-5}$ 4.64 -9.28
EmbMarker+LL-LFEA 75.08 0.01 -0.70 1.39
AGNews Original 93.74 >0.03 0.72 -1.46
RedAlarm 93.74 >0.09 -2.04 4.07
EmbMarker 93.66 <$10^{-9}$ 12.85 -25.70
EmbMarker+LL-LFEA 91.86 0.005 0.48 -0.96
Enron Spam Original 94.74 >0.03 -0.21 0.42
RedAlarm 94.87 >0.47 -0.50 1.00
EmbMarker 94.78 <$10^{-6}$ 6.17 -12.34
EmbMarker+LL-LFEA 92.40 0.01 0.19 -0.39
B.2 Effectiveness of NSmark  on larger LMs
We further test the watermark effectiveness on larger models, including BERT-large-uncased, [^2] https://huggingface.co/google-bert/bert-large-uncased RoBERTa-large, [^3] https://huggingface.co/FacebookAI/roberta-large GPT-2, [^4] https://huggingface.co/openai-community/gpt2 and Llama-2-7B. [^5] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf Experimental results in Table 10 show the effectiveness of NSmarkacross different size of models.
Table 10: Effectiveness of watermark on larger LMs. Both WER and NSMD on larger models are similar to the main results, demonstrating the scalability of NSmark.
Metric $f_{wm}$ $f_{clean}$
BERT-large RoBERTa-large GPT-2 Llama-2-7B BERT-large RoBERTa-large GPT-2 Llama-2-7B
WER $1.00$ $1.00$ $1.00$ $1.00$ $0.00$ $0.03$ $0.00$ $0.00$
NSMD $2.08\times 10^{-8}$ $1.99\times 10^{-6}$ $3.29\times 10^{-6}$ $2.97\times 10^{-6}$ $72.59$ $71.22$ $80.33$ $82.94$
B.3 Defense against LL-LFEA+finetuning
After applying LL-LFEA, the attacker may add a network to $f_{LL-LFEA}$ and fine-tune it for downstream tasks. We hope the model after the LL-LFEA+fine-tuning attack can still maintain the watermark. Table 11 shows results on different LMs and different downstream tasks are not exactly the same. WER of different models has decreased significantly to varying degrees. Most NSMDs are still below the threshold, but RoBERTa and DeBERTa change more on SST-2, which is generally consistent with that of fine-tuning without LL-LFEA attack (Table 4). Through ACC, we can find that LL-LFEA attack does not affect the performance of the model on the original task.
Table 11: Impact of LL-LFEA+ fine-tuning attack on watermark. (i) WER increases to varying degrees; (ii) NSMD are still below the threshold; (iii) ACC does not decrease significantly.
@@ B.3 Defense against LL-LFEA+finetuning
B.4 Robustness against pruning
Table 12 shows results that the watermarked model is pruned and finetuned on the SST-5 dataset. The accuracy of the model only changes slightly.
Table 12: Results of different pruning rates on ACC of watermarked model.
Pruning Rate 0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
ACC(%) 52.62 52.36 52.13 51.99 52.35 51.62 51.94 51.71 50.81 47.23
B.5 Robustness against Fine-pruning
Table 13: Impact of fine-pruning attack on watermark. Generally the results are similar to results of pruning attack in Figure 7 and Table 12.
@@ B.5 Robustness against Fine-pruning
B.6 Robustness against paraphrasing attack
As proposed in (Shetty et al., 2024), paraphrasing attack can bypass many watermark schemes. Thus, we study the robustness of proposed NSmark  against paraphrasing attack. In principle, paraphrasing attack changes the input text and the hidden state of the LM output, which may affect the extraction of WER. However, this does not change the semantic space to which the output matrix belongs, so the output space used to calculate NSMD will not change significantly. Referring to the original paper, we use DIPPER (Krishna et al., 2024) to paraphrase $P=3$ on first $5000$ lines of WikiText-2 and experiment on them. First watermarked model generates embedding, and the averaged embedding of multiple paraphrases are used to train the surrogate model. Results in Table 14 confirms our analysis, and our NSMD indicator is still below the threshold and could be used to verify the watermark.
Table 14: Results of paraphrasing attack on watermarked model. Referring to the original paper, DIPPER (Krishna et al., 2024) is used to paraphrase $P=3$ samples on first $5000$ lines of WikiText-2. Then the samples are used generate average embedding and for training surrogate model.
Model BERT RoBERTa DeBERTa XLNet
WER 0.00 0.00 0.00 0.00
NSMD 15.56 15.11 14.67 13.13
B.7 Robustness against multi-time LL-LFEA attack
Since LL-LFEA has little damage on the model performance, the attacker may try to further destroy the watermark through multiple LL-LFEA attacks. In principle, multiple LL-LFEA will only decrease WER but will not affect NSMD. Table 15 shows the results consistent with our analysis.
Table 15: The results of multi-time LL-LFEA attack on watermark performance.
Number of LL-LFEA 0 1 2 3
WER 1.00 0.27 0.00 0.00
NSMD $2.94\times 10^{-6}$ 0.06 0.04 0.05
B.8 Computational cost analysis
Computation cost of NSmark  basically aligns with existing schemes. Concretely, the cost involves two segments, including model-related and model-unrelated. For model-related computation, such as training of extractor (a three-layer MLP) and assistance of reference model (copy of original LM, only for inference), they are only used in watermark embedding, which is executed only once for each model, and this process is performed in the model training side with a lot of computing power. For model-unrelated computation, it involves a lot of mathematics and cryptography mechanisms, including signature algorithms and hash algorithms. The forward calculation of these cryptographic algorithms consumes have almost no cost on current computing devices, and attackers who want to steal watermarks cannot crack them (computationally unrealistic costs). In contrast, other existing schemes such as WET (Shetty et al., 2024) need to generate a large amount of data through ChatGPT (cost of more that \$100) or DIPPER (Krishna et al., 2024) (11B model), which takes much more time, computation, and money than NSmark. We test the cost of NSmark  on a single Nvidia RTX 3090. The model used is BERT, and other settings are same as Section 4.1. Table 16 shows the results of the time cost (s: second, h: hour). Model_emb means the total time of watermark model training, and Original_model_train means the time of training model without watermark. It can be concluded that NSmark  is pratical to real-world applications.
Table 16: The time cost of NSmark. All results are tested on a single Nvidia RTX 3090. The model used is BERT, and other settings are same as Section 4.1.s denotes second and h denotes hour. Model_emb means the total time of watermark model training, and Original_model_train means the time of training model without watermark.
Process Sign Encode Select SM DSM Gen_Q Cal_NS Model_emb Original_model_train WM_verify
Time 10-4s 10-5s 10-3s 0.5s 0.03s 0.27s 0.83s 4h 3h 37s
B.9 Feature Visualization
To demonstrate the effectiveness of our scheme, we use t-SNE to visualize the feature distribution of the watermarking model. As shown in Figure 9, the input with trigger and the input without trigger can be well separated in the output of LM and $E$, whether for $f_{wm}$ or fine-tuned $F_{wm}$.
Figure 9: The t-SNE visualization of output feature vectors of watermarked models. (i) Left column: $f_{wm}$ on WikiText; (ii) Middle column: $F_{wm}$ on SST-2; (iii) Right column: $F_{wm}$ on SST-5.